# reference.bib

### v1

@@ I Introduction
In recent years, smart agriculture has emerged as a transformative approach to increase farming efficiency, reduce costs, and maintain environmental sustainability [ghazal2024computer, li2022reinforcement]. By incorporating cutting-edge technologies such as the Internet of Things (IoT), artificial intelligence (AI), and advanced data analytics, smart agriculture offers improved monitoring and decision-making across the entire agricultural supply chain. One crucial component of many smart farming solutions is object detection, which enables systems to identify crops, weeds, pests, and machinery in real-time through sensor networks or robotic platforms [akbar2024comprehensive]. This real-time detection capability is instrumental in tasks like early pest detection [lippi2022data] and precise harvesting [zhang2020multi], all of which can significantly enhance yield and utilization of resource [li2024foundation].
@@
Traditional object detection techniques, typically based on convolutional neural networks (CNNs) like Faster R-CNN [zhang2024deep] or YOLO [li2021q], have demonstrated considerable success in classifying and localizing objects such as crops, pests, and agricultural machinery [9754540]. These models typically rely on visual features—such as color, texture, and shape—and perform effectively when trained on sufficiently large and representative datasets [lin2024enhancing]. However, their performance can be limited in complex agricultural environments where visually similar objects might have different contextual meanings (e.g., healthy leaves versus early-stage blight) and annotations are costly to obtain. By contrast, vision-language models integrate textual and visual information, enabling them to capture more semantic details than purely visual models [liu2024grounding, tran2023ai, li2023knowledge]. In agricultural tasks, for example, vision-language models can interpret textual descriptions of symptoms or growth stages, providing richer insights and more accurate identification of diseases, nutrient deficiencies, or weed species [8961482]. Consequently, these models offer a powerful alternative, enhancing interpretability and reducing misclassification by combining linguistic cues with visual signals [zhang2024vision]. While traditional object detection methods remain integral for many baseline tasks, the added semantic understanding of vision-language approaches increasingly positions them as a more comprehensive solution for intelligent farming systems [zhu2024harnessing, liu2024multimodal].
@@
Despite the effectiveness of both traditional and vision-language object detection methods, large-scale data collection is often required to ensure robust performance across diverse agricultural environments [li2024secure]. However, assembling such extensive datasets is challenging. In particular, those containing high-resolution farmland images and potentially identifiable individuals raise significant privacy concerns. These concerns include meeting requirements set forth by regulations such as the General Data Protection Regulation (GDPR) [gupta2020security]. Federated learning (FL) has emerged as a promising approach to address these issues by training a global model across multiple decentralized clients without requiring the direct exchange of raw data. Instead, each client conducts local training and only shares updated parameters, preserving data confidentiality and reducing the risk of unauthorized data exposure [li2020federated, gaur2024impact].
@@
However, standard federated learning solutions introduce another obstacle: communication overhead [al2021reducing]. In a typical FL setup, models—often quite large in object detection or vision-language tasks—must be frequently synchronized between clients and a central server. This repeated transmission of model weights can be bandwidth-intensive, especially in agricultural environments where network connections may be limited or costly [zhang2024vision].For example, a typical object detection model like YOLOv3 includes approximately 62 million trainable parameters, totaling 1,984 Mb. At a transmission speed of 100 Mbps, it takes around 19 seconds for each client to upload the model to the server in each round of federated learning. As a result, practitioners often face a trade-off between privacy protection and practical efficiency; they might reduce synchronization frequency to save bandwidth at the risk of slower convergence or degraded accuracy.
@@
To address the aforementioned challenges, we propose a Vision-Language model-based Lightweight Federated Learning framework (VLLFL) that integrates the semantic strengths of multimodal approaches with the privacy-preserving benefits of federated learning, while simultaneously minimizing communication overhead. An example structure of VLLFL in smart agriculture is presented in Fig. 1. Inspired by the design philosophy presented in [qiu2023text], which demonstrated the feasibility of training prompt generators in a federated manner for classification tasks, our approach focuses on learning a compact prompt generator that interacts with the model’s text encoder. This strategy significantly reduces the number of parameters that must be exchanged, thereby lowering bandwidth consumption and accelerating convergence. In this way, VLLFL preserves the contextual understanding of vision-language models for object detection, ensures stringent data privacy through federated learning, and enables an efficient model update process suitable for large-scale agricultural deployments. In our experiments, VLLFL significantly enhanced the object detection accuracy of the VLM for fruit detection across multiple classes and clients. Specifically, VLLFL achieved an average improvement in mean average precision (mAP) from 9.59% to 24.12% while reduce communication overhead by 99.3%, thereby demonstrating VLLFL’s superior efficiency in deploying vision-language models via federated learning.
@@ II-A Vision Language Model
Vision-language models have become increasingly popular for their ability to jointly process textual and visual content, unlocking more context-aware performance than single-modality architectures [li2024vision]. By integrating linguistic cues—such as object labels or domain-specific descriptions—with visual features, these models capture semantic nuances often overlooked by traditional object detection approaches, which typically rely on attributes like color, texture, and shape. For example, while a conventional detector might classify an object simply as an “apple” based on visual characteristics, a vision-language model can further distinguish between varieties such as “green apple” and “red apple” if such labels are provided. This cross-modal capability not only enhances detection precision but also facilitates a deeper interpretation of objects in real-world environments—particularly valuable in agriculture, where visually similar varieties often demand fine-grained differentiation [ghosh2024exploring].
In recent years, numerous widely adopted vision-language models have demonstrated their versatility across various domains, including agriculture. Early approaches, such as the two-tower architecture proposed by [kiros2014unifying], aligned separate image and text embeddings within a shared space, enabling basic cross-modal retrieval. More advanced transformer-based models—including ViLBERT [lu2019vilbert], LXMERT [tan2019lxmert], and UNITER [chen2020uniter]—introduced multi-layer cross-attention mechanisms, significantly improving joint image-text representations. Trained on large-scale datasets, these models achieved strong performance in tasks such as visual question answering and referring expression comprehension. Subsequently, CLIP [radford2021learning] achieved impressive zero-shot classification by learning generalized representations from large-scale image-text corpora, while DALL·E [ramesh2021zero] expanded the frontier by generating images from textual prompts. In agricultural contexts, such models offer promising capabilities for detecting pests, identifying crop diseases, or generating descriptive captions of field conditions—providing farmers with detailed, context-rich insights to support informed decision-making.
Among vision-language paradigms, GroundingDINO [liu2024grounding] stands out for its integration of object detection with language grounding. It extends DINO’s detection-oriented architecture by incorporating mechanisms that align textual prompts or labels with corresponding image regions. This design excels at generating high-quality bounding boxes guided by semantic cues—a capability particularly valuable in agricultural applications where fine-grained distinctions (e.g., between plant varieties or infection stages) are essential. As a result, we adopt GroundingDINO as the backbone of our proposed framework, leveraging its strong detection performance and linguistic grounding capabilities to advance precision agriculture.
@@ II-B Federated-Learning with Vision-Language Models
Federated learning (FL) has emerged as a powerful paradigm for enabling collaborative model training without directly sharing raw data. While it has been successfully applied to a range of machine learning tasks—such as image classification and natural language processing—its adoption in vision-language contexts is comparatively recent. The fundamental idea behind FL is to train a global model by aggregating locally computed updates (e.g., gradients or model weights) from multiple clients. This setup preserves data privacy and often complies more readily with regulations (e.g., GDPR) because the sensitive information remains on local devices or servers [li2020federated]. However, deploying large-scale vision-language models within an FL framework introduces unique challenges related to communication overhead, data heterogeneity, and model complexity [che2023multimodal].
@@
Vision-language models typically consist of high-dimensional parameters due to their attention-based architectures and large-scale pretraining objectives [lu2019vilbert]. For instance, the GroundingDINO model discussed earlier contains approximately 172 million trainable parameters, in total of 656 MB. In federated learning settings, training such large models can result in substantial communication overhead, as model updates must be exchanged between clients and a central server. This issue is especially critical in domains like agriculture and healthcare, where network bandwidth is often limited and connectivity may be intermittent. Existing strategies to mitigate overhead include parameter-efficient fine-tuning methods—such as prompt tuning, adapter layers, or low-rank factorization—and gradient compression techniques that reduce the payload of each synchronization step. These approaches help make federated learning of vision-language models more feasible and cost-effective.
@@
On the data side, federated learning for vision-language tasks often involves heterogeneous and domain-specific inputs [nguyen2024flora]. In agriculture, for instance, local nodes may have imagery that corresponds to different crop varieties, growth conditions, or climate zones, alongside textual annotations that vary widely in specificity and language . This diversity can complicate the alignment of textual and visual modalities across clients, often necessitating advanced domain adaptation and continual learning strategies to ensure robust global model performance. Nonetheless, when effectively implemented, FL confers significant advantages for vision-language applications: it preserves data confidentiality, caters to multi-farm or multi-region collaborations, and leverages distributed resources for large-scale model training.
@@
Recent frameworks have shown promise in bridging vision-language models and federated learning. Some efforts focus on specialized tasks like caption generation or referring expression comprehension, relying on partial updates (e.g., prompt generators) instead of full model weights to conserve bandwidth [yan2024lightweight]. Others explore knowledge distillation methods, where a centrally maintained teacher model guides local models, further reducing synchronization burden [pan2024survey]. As advances in efficient training techniques and communication protocols continue, federated learning with vision-language models is poised to open up collaborative opportunities in industries where data sensitivity and distribution constraints are paramount, from smart agriculture to healthcare and beyond.
Figure 2: The architecture of VLM backbone in VLLFL based on GroundingDINO[liu2024grounding]
@@ II-C Vision-Language Model Prompt Learning
Prompt learning has emerged as an effective way to adapt large vision-language models without fine-tuning the entirety of their parameters [bu2024efficient]. Rather than modifying all layers, a small number of learnable “prompt tokens” or “prompt vectors” is inserted into the model, guiding its attention or output distribution toward a target task. This approach outperforms rigid, hand-crafted prompts (e.g., “a photo of a [CLASS]”) by dynamically optimizing textual cues, thus enabling more nuanced adaptations. For agriculture, prompt learning is particularly appealing because complex domain-specific vocabulary—covering different crop varieties or pest types—can be integrated without excessively retraining the underlying model [li2024foundation].
Recently, researchers have proposed a range of prompt learning methods that build on this principle. Context Optimization (CoOp) and Conditional Context Optimization (Co-CoOp) [zhou2022conditional] introduce trainable prompts to large-scale vision-language models like CLIP [radford2021learning], achieving strong results in zero-shot or few-shot classification settings. Efforts like FedTPG [qiu2024federated] extend prompt learning to federated learning scenarios, focusing on a lightweight prompt generator shared across multiple clients. By exchanging minimal parameter updates, FedTPG reduces communication overhead compared to conventional federated approaches. At the same time, frameworks such as PromptDet [feng2022promptdet] explore prompt-based strategies for object detection; however, these methods do not currently operate in a federated context, often assuming centralized data availability for training detection models.
To date, there has been no dedicated prompt learning solution tailored specifically for federated object detection tasks. This gap is particularly notable in smart agriculture, where privacy constraints and distributed data ownership are paramount. Inspired by FedTPG’s success in federated classification with prompt learning, we propose VLLFL—a federated learning framework that leverages prompt tuning for object detection in real-world agricultural environments [qiu2023text]. By focusing on training compact prompts rather than entire model weights, VLLFL seeks to preserve data privacy, minimize communication overhead, and deliver high-performing detection capabilities for tasks such as identifying crop diseases, detecting wildlife threats, or monitoring farm machinery.
@@ III-A Vision-Language Model
Vision-language models excel in zero-shot and open-set recognition by leveraging text-based cues to detect previously unseen object categories, even without prior training examples [li2024foundation]. A typical VLM architecture consists of a text encoder and an image encoder, which extract features from textual and visual inputs, respectively, and project them into a shared vector space. Predictions are then made by matching the encoded image features with their corresponding text features. This shared visual-language representation enhances robustness to domain shifts, enabling strong generalization across diverse styles and contexts. Moreover, VLMs can be rapidly adapted to a wide range of downstream tasks with minimal labeled data.
In our proposed VLLFL framework, the foundation model employed for object detection is GroundingDINO [liu2024grounding]. GroundingDINO is a popular open-set detector built upon the DETR-like architecture DINO [zhang2022dino], which integrates end-to-end transformer-based detection mechanisms. The overall workflow of GroundingDINO is illustrated in Fig. 2. Initially, the model extracts visual and textual features using an image encoder (e.g., the Swin Transformer [liu2021swin]) and a text encoder (e.g., BERT [devlin2019bert]), respectively. These extracted features are then passed to a feature enhancer network, which fuses cross-modal information to enable comprehensive integration of textual and visual cues.
@@ III-C Federated-learning Model Aggregation
Algorithm 1 VLLFL Algorithm [qiu2024federated]
@@
7:         Send the current global model $\theta^{r}$ to client $i$.
8:         Receive the locally updated model parameters $\theta_{i}^{r+1}$ from Local Client $i$.
9:     end for
10:     Aggregate the model parameters:
11:     $\theta^{r+1}\leftarrow\frac{1}{|S^{r}|}\sum_{i\in S^{r}}\theta_{i}^{r+1}$
@@
19:     Concatenate the prompt vectors with the class embedding

### v2

# VLLFL: A Vision-Language Model Based Lightweight Federated Learning Framework for Smart Agriculture

@@ I Introduction
In recent years, smart agriculture has emerged as a transformative approach to increase farming efficiency, reduce costs, and maintain environmental sustainability [1, 2]. By incorporating cutting-edge technologies such as the Internet of Things (IoT), artificial intelligence (AI), and advanced data analytics, smart agriculture offers improved monitoring and decision-making across the entire agricultural supply chain [3]. One crucial component of many smart farming solutions is object detection, which enables systems to identify crops, weeds, pests, and machinery in real-time through sensor networks or robotic platforms [4]. This real-time detection capability is instrumental in tasks like early pest detection [5] and precise harvesting [6], all of which can significantly enhance yield and utilization of resource [7].
@@
Traditional object detection techniques, typically based on convolutional neural networks (CNNs) like Faster R-CNN [8] or YOLO [9], have demonstrated considerable success in classifying and localizing objects.[10]. These models typically rely on visual features—such as color, texture, and shape—and perform effectively when trained on sufficiently large and representative datasets [11]. However, their performance can be limited in complex agricultural environments where visually similar objects might have different contextual meanings (e.g., healthy leaves versus early-stage blight) and annotations are costly to obtain. By contrast, vision-language models integrate textual and visual information, enabling them to capture more semantic details than purely visual models [12, 13, 14]. In agricultural tasks, for example, vision-language models can interpret textual descriptions of symptoms or growth stages, providing richer insights and more accurate identification of diseases, nutrient deficiencies, or weed species [15]. Consequently, these models offer a powerful alternative, enhancing interpretability and reducing misclassification by combining linguistic cues with visual signals [16]. While traditional object detection methods remain integral for many baseline tasks, the added semantic understanding of vision-language approaches increasingly positions them as a more comprehensive solution for intelligent farming systems [17, 18].
@@
Despite the effectiveness of both traditional and vision-language object detection methods, large-scale data collection is often required to ensure robust performance across diverse agricultural environments [19]. However, assembling such extensive datasets is challenging. In particular, those containing high-resolution farmland images and potentially identifiable individuals raise significant privacy concerns [20]. These concerns include meeting requirements set forth by regulations such as the General Data Protection Regulation (GDPR) [21]. Federated learning (FL) has emerged as a promising approach to address these issues by training a global model across multiple decentralized clients without requiring the direct exchange of raw data. Instead, each client conducts local training and only shares updated parameters, preserving data confidentiality and reducing the risk of unauthorized data exposure [22, 23].
@@
However, standard federated learning solutions introduce another obstacle: communication overhead [24]. In a typical FL setup, models—often quite large in object detection or vision-language tasks—must be frequently synchronized between clients and a central server. This repeated transmission of model weights can be bandwidth-intensive, especially in agricultural environments where network connections may be limited or costly [16].For example, a typical object detection model like YOLOv3 includes approximately 62 million trainable parameters, totaling 1,984 Mb. At a transmission speed of 100 Mbps, it takes around 19 seconds for each client to upload the model to the server in each round of federated learning. As a result, practitioners often face a trade-off between privacy protection and practical efficiency; they might reduce synchronization frequency to save bandwidth at the risk of slower convergence or degraded accuracy.
@@
To address the aforementioned challenges, we propose a Vision-Language model-based Lightweight Federated Learning framework (VLLFL) that integrates the semantic strengths of multimodal approaches with the privacy-preserving benefits of federated learning, while simultaneously minimizing communication overhead. An example structure of VLLFL in smart agriculture is presented in Fig. 1. Inspired by the design philosophy presented in [25], which demonstrated the feasibility of training prompt generators in a federated manner for classification tasks, our approach focuses on learning a compact prompt generator that interacts with the model’s text encoder. This strategy significantly reduces the number of parameters that must be exchanged, thereby lowering bandwidth consumption and accelerating convergence. In this way, VLLFL preserves the contextual understanding of vision-language models for object detection, ensures stringent data privacy through federated learning, and enables an efficient model update process suitable for large-scale agricultural deployments. In our experiments, VLLFL significantly enhanced the object detection accuracy of the VLM for fruit detection across multiple classes and clients. Specifically, VLLFL achieved an average improvement in mean average precision (mAP) from 9.59% to 24.12% while reduce communication overhead by 99.3%, thereby demonstrating VLLFL’s superior efficiency in deploying vision-language models via federated learning.
@@ II-A Vision Language Model
Vision-language models have become increasingly popular for their ability to jointly process textual and visual content, unlocking more context-aware performance than single-modality architectures [26]. By integrating linguistic cues—such as object labels or domain-specific descriptions—with visual features, these models capture semantic nuances often overlooked by traditional object detection approaches, which typically rely on attributes like color, texture, and shape. For example, while a conventional detector might classify an object simply as an “apple” based on visual characteristics, a vision-language model can further distinguish between varieties such as “green apple” and “red apple” if such labels are provided. This cross-modal capability not only enhances detection precision but also facilitates a deeper interpretation of objects in real-world environments—particularly valuable in agriculture, where visually similar varieties often demand fine-grained differentiation [27].
In recent years, numerous widely adopted vision-language models have demonstrated their versatility across various domains, including agriculture. Early approaches, such as the two-tower architecture proposed by [28], aligned separate image and text embeddings within a shared space, enabling basic cross-modal retrieval. More advanced transformer-based models—including ViLBERT [29], LXMERT [30], and UNITER [31]—introduced multi-layer cross-attention mechanisms, significantly improving joint image-text representations. Trained on large-scale datasets, these models achieved strong performance in tasks such as visual question answering and referring expression comprehension. Subsequently, CLIP [32] achieved impressive zero-shot classification by learning generalized representations from large-scale image-text corpora, while DALL·E [33] expanded the frontier by generating images from textual prompts. In agricultural contexts, such models offer promising capabilities for detecting pests, identifying crop diseases, or generating descriptive captions of field conditions—providing farmers with detailed, context-rich insights to support informed decision-making.
Among vision-language paradigms, GroundingDINO [12] stands out for its integration of object detection with language grounding. It extends DINO’s detection-oriented architecture by incorporating mechanisms that align textual prompts or labels with corresponding image regions. This design excels at generating high-quality bounding boxes guided by semantic cues—a capability particularly valuable in agricultural applications where fine-grained distinctions (e.g., between plant varieties or infection stages) are essential. As a result, we adopt GroundingDINO as the backbone of our proposed framework, leveraging its strong detection performance and linguistic grounding capabilities to advance precision agriculture.
@@ II-B Federated-Learning with Vision-Language Models
Federated learning (FL) has emerged as a powerful paradigm for enabling collaborative model training without directly sharing raw data. While it has been successfully applied to a range of machine learning tasks—such as image classification and natural language processing—its adoption in vision-language contexts is comparatively recent. The fundamental idea behind FL is to train a global model by aggregating locally computed updates (e.g., gradients or model weights) from multiple clients. This setup preserves data privacy and often complies more readily with regulations (e.g., GDPR) because the sensitive information remains on local devices or servers [22]. However, deploying large-scale vision-language models within an FL framework introduces unique challenges related to communication overhead, data heterogeneity, and model complexity [34].
@@
Vision-language models typically consist of high-dimensional parameters due to their attention-based architectures and large-scale pretraining objectives [29]. For instance, the GroundingDINO model discussed earlier contains approximately 172 million trainable parameters, in total of 656 MB. In federated learning settings, training such large models can result in substantial communication overhead, as model updates must be exchanged between clients and a central server. This issue is especially critical in domains like agriculture and healthcare, where network bandwidth is often limited and connectivity may be intermittent. Existing strategies to mitigate overhead include parameter-efficient fine-tuning methods—such as prompt tuning, adapter layers, or low-rank factorization—and gradient compression techniques that reduce the payload of each synchronization step. These approaches help make federated learning of vision-language models more feasible and cost-effective.
@@
On the data side, federated learning for vision-language tasks often involves heterogeneous and domain-specific inputs [35]. In agriculture, for instance, local nodes may have imagery that corresponds to different crop varieties, growth conditions, or climate zones, alongside textual annotations that vary widely in specificity and language . This diversity can complicate the alignment of textual and visual modalities across clients, often necessitating advanced domain adaptation and continual learning strategies to ensure robust global model performance. Nonetheless, when effectively implemented, FL confers significant advantages for vision-language applications: it preserves data confidentiality, caters to multi-farm or multi-region collaborations, and leverages distributed resources for large-scale model training.
@@
Recent frameworks have shown promise in bridging vision-language models and federated learning. Some efforts focus on specialized tasks like caption generation or referring expression comprehension, relying on partial updates (e.g., prompt generators) instead of full model weights to conserve bandwidth [36]. Others explore knowledge distillation methods, where a centrally maintained teacher model guides local models, further reducing synchronization burden [37]. As advances in efficient training techniques and communication protocols continue, federated learning with vision-language models is poised to open up collaborative opportunities in industries where data sensitivity and distribution constraints are paramount, from smart agriculture to healthcare and beyond.
Figure 2: The architecture of VLM backbone in VLLFL based on GroundingDINO[12]
@@ II-C Vision-Language Model Prompt Learning
Prompt learning has emerged as an effective way to adapt large vision-language models without fine-tuning the entirety of their parameters [38]. Rather than modifying all layers, a small number of learnable “prompt tokens” or “prompt vectors” is inserted into the model, guiding its attention or output distribution toward a target task. This approach outperforms rigid, hand-crafted prompts (e.g., “a photo of a [CLASS]”) by dynamically optimizing textual cues, thus enabling more nuanced adaptations. For agriculture, prompt learning is particularly appealing because complex domain-specific vocabulary—covering different crop varieties or pest types—can be integrated without excessively retraining the underlying model [7].
Recently, researchers have proposed a range of prompt learning methods that build on this principle. Context Optimization (CoOp) and Conditional Context Optimization (Co-CoOp) [39] introduce trainable prompts to large-scale vision-language models like CLIP [32], achieving strong results in zero-shot or few-shot classification settings. Efforts like FedTPG [40] extend prompt learning to federated learning scenarios, focusing on a lightweight prompt generator shared across multiple clients. By exchanging minimal parameter updates, FedTPG reduces communication overhead compared to conventional federated approaches. At the same time, frameworks such as PromptDet [41] explore prompt-based strategies for object detection; however, these methods do not currently operate in a federated context, often assuming centralized data availability for training detection models.
To date, there has been no dedicated prompt learning solution tailored specifically for federated object detection tasks. This gap is particularly notable in smart agriculture, where privacy constraints and distributed data ownership are paramount. Inspired by FedTPG’s success in federated classification with prompt learning, we propose VLLFL—a federated learning framework that leverages prompt tuning for object detection in real-world agricultural environments [25]. By focusing on training compact prompts rather than entire model weights, VLLFL seeks to preserve data privacy, minimize communication overhead, and deliver high-performing detection capabilities for tasks such as identifying crop diseases, detecting wildlife threats, or monitoring farm machinery.
@@ III-A Vision-Language Model
Vision-language models excel in zero-shot and open-set recognition by leveraging text-based cues to detect previously unseen object categories, even without prior training examples [7]. A typical VLM architecture consists of a text encoder and an image encoder, which extract features from textual and visual inputs, respectively, and project them into a shared vector space. Predictions are then made by matching the encoded image features with their corresponding text features. This shared visual-language representation enhances robustness to domain shifts, enabling strong generalization across diverse styles and contexts. Moreover, VLMs can be rapidly adapted to a wide range of downstream tasks with minimal labeled data.
In our proposed VLLFL framework, the foundation model employed for object detection is GroundingDINO [12]. GroundingDINO is a popular open-set detector built upon the DETR-like architecture DINO [42], which integrates end-to-end transformer-based detection mechanisms. The overall workflow of GroundingDINO is illustrated in Fig. 2. Initially, the model extracts visual and textual features using an image encoder (e.g., the Swin Transformer [43]) and a text encoder (e.g., BERT [44]), respectively. These extracted features are then passed to a feature enhancer network, which fuses cross-modal information to enable comprehensive integration of textual and visual cues.
@@ III-C Federated-learning Model Aggregation
Algorithm 1 VLLFL Algorithm [40]
@@
7:   Receive the local model parameters $\theta_{i}^{r+1}$ from Client ii.
8:  end for
9:  Aggregate the model parameters:
10:  $\theta^{r+1}\leftarrow\frac{1}{|S^{r}|}\sum_{i\in S^{r}}\theta_{i}^{r+1}$
11:  Distribute the current global model $\theta_{i}^{r+1}$ to all selected clients.
@@
19:  Concatenate the prompt vectors with class embedding